# Quantum field theory treatment of oscillations of Dirac neutrinos in external fields

### v1

@@ 1 Introduction
The key issue in this approach is using the propagators of neutrino mass eigenstates in the calculation of the matrix elements. If neutrinos propagate between a source and a detector in vacuum, the propagators have a well-known form. It allows one to calculate the transition probability for neutrino oscillations in vacuum. The generalization of this formalism for the description of neutrino oscillations in external fields is nontrivial since, as we shall see shortly, neutrino mass eigenstates are mixed in these situations. One can apply the QFT formalism in question straightforwardly only in a particular case when an external field is diagonal in the mass eigenstates basis (see, e.g., Ref. [16]).
@@
The present work is organized in the following way. After recalling the basic issues of neutrino masses and mixing in Sec. 2, we describe how QFT is applied to neutrino flavor oscillations in external fields in Sec. 3. Then, in Sec. 4, we study the situation of neutrinos propagating in background matter. We remind how neutrinos can interact with matter within the standard model in Sec. 4.1. Then, we derive the dressed propagators of Dirac neutrinos interacting with background matter in Sec. 4.2. The matrix element and the transition probability for oscillations of Dirac neutrinos in matter are obtained in Sec. 4.3.
The impact of a magnetic field on the spin-flavor precession of Dirac neutrinos in frames of QFT is considered in Sec. 5. We briefly recall the basics of the electrodynamics of Dirac neutrinos in Sec. 5.1. The QM description of the spin-flavor precession of Dirac neutrinos having a transition magnetic moment is described in Sec. 5.2. The dressed propagators of Dirac neutrinos in a magnetic field are derived in Sec. 5.3. The matrix element and the corresponding probabilities of the spin-flavor precession of Dirac neutrinos in frames of QFT are obtained in Sec. 5.4. Finally, we conclude in Sec. 6.
The diagonal propagators of Dirac neutrinos in matter and in vacuum are derived in Appendices A and C. We provide some details for the computation of the matrix elements in Appendix B.
@@ 2 Neutrino masses and mixing
In this work, we adopt a specific situation when only a Dirac mass term is present, i.e. the corresponding mass term in the Lagrangian reads
@@
where $\psi_{a}$ are Dirac particles, i.e. $\psi_{a}^{c}\neq\psi_{a}$, having masses $m_{a}$, and $\gamma^{\mu}=(\gamma^{0},\bm{\gamma})$ are the Dirac matrices.
@@ 3 QFT treatment of neutrino oscillations
$\left\langle l_{\alpha}\left|T\left\{j_{\mu}^{\dagger}(x)j^{\nu}(y)\right\}\right|l_{\beta}\right\rangle=$ $e^{\mathrm{i}p_{\alpha}x-\mathrm{i}p_{\beta}y}\bar{u}_{\alpha}(p_{\alpha})\gamma_{0}^{\mathrm{L}}$
$\times\sum_{ab}U_{\alpha a}U_{\beta b}^{*}\left\langle 0\left|T\left\{\psi_{a}(x)\bar{\psi}_{b}(y)\right\}\right|0\right\rangle\gamma_{0}^{\mathrm{L}}u_{\beta}(p_{\beta}).$ (3.4)
@@ 4.2 Dressed propagators of Dirac neutrinos in matter
$S_{1}^{-1}-GS_{2}G$ $=\Sigma_{11}^{-1},\quad S_{2}^{-1}-GS_{1}G=\Sigma_{22}^{-1},$ (4.7)
@@
$S_{a}^{-1}$ $=-\frac{4(p_{0}-E_{a-}-g_{a}/2+\mathrm{i}0)}{(1-\alpha_{a}^{2})(1-\xi^{2})}\left(\begin{array}[]{cc}0&1+\xi(\bm{\sigma}\hat{p})\\ 1+\xi(\bm{\sigma}\hat{p})&0\end{array}\right)(1+\alpha_{a}\gamma^{5}).$ (4.11)
@@
$D_{1}^{-1}$ $=\frac{g(1+\alpha)/2-A(a-b)(1+\xi^{2})+2A\xi(a-b)(\bm{\sigma}\hat{p})}{\left[g(1+\alpha)/2-A(a-b)(1+\xi^{2})\right]^{2}-4A^{2}\xi^{2}(a-b)^{2}},$
$D_{2}^{-1}$ $=\frac{g(1-\alpha)/2-A(a+b)(1+\xi^{2})+2A\xi(a+b)(\bm{\sigma}\hat{p})}{\left[g(1-\alpha)/2-A(a+b)(1+\xi^{2})\right]^{2}-4A^{2}\xi^{2}(a+b)^{2}}.$ (4.15)
@@
$D_{1}^{-1}\to-\frac{g[1-(\bm{\sigma}\hat{p})]}{2[(p_{0}-E_{1-}-g_{1}/2+\mathrm{i}0)(p_{0}-E_{2-}-g_{2}/2+\mathrm{i}0)-g^{2}]},$ (4.16)
@@
$\Sigma_{12}=-\frac{g}{2[(p_{0}-E_{1-}-g_{1}/2+\mathrm{i}0)(p_{0}-E_{2-}-g_{2}/2+\mathrm{i}0)-g^{2}]}\left(\begin{array}[]{cc}0&0\\ 1-(\bm{\sigma}\hat{p})&0\end{array}\right),$ (4.17)
@@
$\Sigma_{11}$ $=-\frac{p_{0}-E_{2-}-g_{2}/2}{2[(p_{0}-E_{1-}-g_{1}/2+\mathrm{i}0)(p_{0}-E_{2-}-g_{2}/2+\mathrm{i}0)-g^{2}]}\left(\begin{array}[]{cc}0&0\\ 1-(\bm{\sigma}\hat{p})&0\end{array}\right),$ (4.20)
$\Sigma_{22}$ $=-\frac{p_{0}-E_{1-}-g_{1}/2}{2[(p_{0}-E_{1-}-g_{1}/2+\mathrm{i}0)(p_{0}-E_{2-}-g_{2}/2+\mathrm{i}0)-g^{2}]}\left(\begin{array}[]{cc}0&0\\ 1-(\bm{\sigma}\hat{p})&0\end{array}\right).$ (4.23)
@@
$S_{a}\to-\frac{1}{2(p_{0}-E_{a-}-g_{a}/2+\mathrm{i}0)}\left(\begin{array}[]{cc}0&0\\ 1-(\bm{\sigma}\hat{p})&0\end{array}\right).$ (4.24)
@@ 4.3 Matrix element and transition probability in matter
To proceed with the calculation of the matrix element in Eq. (3) we choose the coordinate system so that $\mathbf{L}=L\mathbf{e}_{z}$, with $L>0$. Then, for simplicity, we assume the forward scattering approximation for charged leptons. In this case, $u_{e,\mu}^{(\mathrm{L})\mathrm{T}}=(0,\kappa_{e,\mu})$, where $\kappa_{e,\mu}^{\mathrm{T}}=(0,1)$. Thus, both the incoming electron and the outgoing muon are left particles propagating along the $z$-axis. Moreover, we rewrite the energies of left neutrinos as $E_{a-}\approx E_{a}+g_{a}/2$, where $E_{a}=\sqrt{q^{2}+m_{a}^{2}}$ is the energy of a free particle. We also use the cylindrical coordinates $(\rho,z,\phi)$ for the momentum of a virtual neutrino, $\mathbf{q}=\rho\mathbf{e}_{\rho}+z\mathbf{e}_{z}$.
@@
$\mathcal{M}_{e\to\mu}=-\frac{\mathrm{i}\bar{E}_{m}e^{\mathrm{i}\bar{E}_{m}L}}{2\pi L\mathfrak{E}_{m}}\left[\left(\frac{\Delta m^{2}}{4E}+\frac{g_{2}-g_{1}}{2}\right)\sin 2\theta+g\cos 2\theta\right]\sin(\mathfrak{E}_{m}L),$ (4.27)
@@
and $\Delta m^{2}=m_{2}^{2}-m_{1}^{2}>0$.
@@
with help of Eqs. (2.2) and (4.4). Taking into account Eqs. (4.2) and (4.27)-(4.29), one gets the transition probability for flavor oscillations of Dirac neutrinos in matter,
$P_{\nu_{e}\to\nu_{\mu}}\propto$ $|\mathcal{M}_{e\to\mu}|^{2}\propto\frac{\left(\frac{\Delta m^{2}}{4E}\sin 2\theta\right)^{2}}{\left(\frac{\Delta m^{2}}{4E}\sin 2\theta\right)^{2}+\left(\frac{\Delta m^{2}}{4E}\cos 2\theta-\frac{G_{\mathrm{F}}n_{e}}{\sqrt{2}}\right)^{2}}$
$\times\sin^{2}\left(\sqrt{\left(\frac{\Delta m^{2}}{4E}\sin 2\theta\right)^{2}+\left(\frac{\Delta m^{2}}{4E}\cos 2\theta-\frac{G_{\mathrm{F}}n_{e}}{\sqrt{2}}\right)^{2}}L\right).$ (4.30)
Formally, Eq. (4.3) coincides with the prediction of the QM approach for neutrino oscillations in matter [2, 3]. However, in QM, the quantity $E$ is referred to a mean energy of a neutrino beam, $E_{\nu}$, which is a not well defined parameter if one deals with a system of flavor neutrinos. We recall that, in the QFT treatment, $E=(E_{\alpha}+E_{\beta})/2$. Thus, this quantity is related to the energies of charged leptons, which are the observables since these particles correspond to in and out states.
@@ 5.1 Interaction of Dirac neutrinos with an electromagnetic field
where $(U_{a\lambda})=U^{\mathrm{T}}$ is the transposed mixing matrix given in Eq. (2.2).
If only the magnetic field is present, $\mathbf{E}=0$ and $\mathbf{B}\neq 0$, we shall see shortly in Sec. 5.2 that the transitions between neutrinos belonging to different flavors and opposite helicity states are possible provided that these neutrinos have a transition magnetic moment. This process is called the neutrino spin-flavor precession.
@@ 5.2 QM description of the neutrino spin-flavor precession
As in Sec. 4, study two flavor neutrinos, $\nu_{e}$ and $\nu_{\mu}$. As a rule, neutrino oscillations are described in the flavor basis in frames of the QM approach. However, owing to the fact that both the nature of neutrinos and the neutrino magnetic moments are defined for neutrino mass eigenstates, we adopt the mass basis here. The corresponding mass eigenstates are supposed to be Dirac particles. We take that these massive neutrinos have only the transition magnetic moment $\mu$ and interact with a magnetic field $\mathbf{B}=(B,0,0)$ transverse with respect to the neutrino momentum which is along the $z$-axis.
@@ 5.3 Dressed propagators of Dirac neutrinos in magnetic field
$\mathbf{l}$ $=\hat{p}\left(\frac{2\varsigma_{1}(p_{0}-E_{1}+\mathrm{i}0)}{1-\varsigma_{1}^{2}}+\frac{\mu^{2}B^{2}\varsigma_{2}}{2(p_{0}-E_{2}+\mathrm{i}0)}\right)-\frac{\mu^{2}\varsigma_{2}(\mathbf{B}\hat{p})\mathbf{B}}{(p_{0}-E_{2}+\mathrm{i}0)}.$ (5.31)
@@
$\Sigma_{11}=\frac{(p_{0}-E_{2})(\gamma^{0}-\bm{\gamma}\hat{p})}{2[(p_{0}-E_{1}+\mathrm{i}0)(p_{0}-E_{2}+\mathrm{i}0)-\mu^{2}B^{2}+\mu^{2}(\mathbf{B}\hat{p})^{2}]}.$ (5.32)
@@
$\Sigma_{22}=\frac{(p_{0}-E_{1})(\gamma^{0}-\bm{\gamma}\hat{p})}{2[(p_{0}-E_{1}+\mathrm{i}0)(p_{0}-E_{2}+\mathrm{i}0)-\mu^{2}B^{2}+\mu^{2}(\mathbf{B}\hat{p})^{2}]}.$ (5.33)
@@
The denominators of $\Sigma_{aa}$ in Eqs. (5.32) and (5.33) contain the term $\propto\mu^{2}(\mathbf{B}\cdot\hat{p})^{2}$. Analogous terms are present in the dressed propagators of Majorana neutrinos in a magnetic field, which were obtained in Ref. [18]. In the QM description of the spin-flavor precession in Sec. 5.2, the magnetic field is taken to be transverse to the neutrino momentum. Thus, $(\mathbf{B}\cdot\hat{p})=0$ since neutrinos are on the mass shell in the QM approach. In QFT treatment, we cannot drop the term $\propto\mu^{2}(\mathbf{B}\cdot\hat{p})^{2}\neq 0$ in Eqs. (5.32) and (5.33) since $\mathbf{p}$ is the momentum of a virtual neutrino. In fact, $\mathbf{p}$ can be arbitrary since we integrate over it in Eq. (3). Note that the quantity in question, $\propto\mu^{2}(\mathbf{B}\cdot\hat{p})^{2}$, is an essentially quantum term in the propagator which arises from the summation of the infinite number of the Feynman diagrams in Fig. 2 Shortly in Sec. 5.4, we shall compute the contributions to the matrix element and the transition probability originating from this term.

### v2

@@ 1 Introduction
The key issue in this approach is using the propagators of neutrino mass eigenstates in the calculation of the matrix elements. If neutrinos propagate between a source and a detector in vacuum, the propagator has a well-known form. It allows one to calculate the transition probability for neutrino oscillations in vacuum. The generalization of this formalism for the description of neutrino oscillations in external fields is nontrivial since, as we shall see shortly, neutrino mass eigenstates are mixed in these situations. One can apply the QFT formalism in question straightforwardly only in a particular case when an external field is diagonal in the mass eigenstates basis (see, e.g., Ref. [16]).
@@
The present work is organized in the following way. After recalling the basic issues of neutrino masses and mixing in Sec. 2, we describe how QFT is applied to neutrino flavor oscillations in external fields in Sec. 3.
Then, in Sec. 4, we study the situation of neutrinos propagating in background matter. We remind how neutrinos can interact with matter within the standard model in Sec. 4.1. Then, we derive the dressed propagators of Dirac neutrinos interacting with background matter in Sec. 4.2. The matrix element and the transition probability for oscillations of Dirac neutrinos in matter are obtained in Sec. 4.3.
The impact of a magnetic field on the spin-flavor precession of Dirac neutrinos in frames of QFT is considered in Sec. 5. We briefly recall the basics of the electrodynamics of Dirac neutrinos in Sec. 5.1. The QM description of the spin-flavor precession of Dirac neutrinos having a transition magnetic moment is described in Sec. 5.2. The dressed propagators of Dirac neutrinos in a magnetic field are derived in Sec. 5.3. The matrix element and the corresponding probabilities of the spin-flavor precession of Dirac neutrinos in frames of QFT are obtained in Sec. 5.4.
We consider some general problems which arise in the application of QFT to neutrino oscillations in Sec. 6. Finally, we conclude in Sec. 7. The diagonal propagators of Dirac neutrinos in matter and in vacuum are derived in Appendices A and C. We provide some details for the computation of the matrix elements in Appendix B.
@@ 2 Neutrino masses and mixing
In this work, we adopt a specific situation when only a Dirac mass term is present, i.e. the corresponding term in the Lagrangian reads
@@
where $\psi_{a}$ are the Dirac fields, i.e. $\psi_{a}^{c}\neq\psi_{a}$, having masses $m_{a}$, and $\gamma^{\mu}=(\gamma^{0},\bm{\gamma})$ are the Dirac matrices.
@@ 3 QFT treatment of neutrino oscillations
$\left\langle l_{\alpha}\left|T\left\{j_{\mu}^{\dagger}(x)j_{\nu}(y)\right\}\right|l_{\beta}\right\rangle=$ $e^{\mathrm{i}p_{\alpha}x-\mathrm{i}p_{\beta}y}\bar{u}_{\alpha}(p_{\alpha})\gamma_{\mu}^{\mathrm{L}}$
$\times\sum_{ab}U_{\alpha a}U_{\beta b}^{*}\left\langle 0\left|T\left\{\psi_{a}(x)\bar{\psi}_{b}(y)\right\}\right|0\right\rangle\gamma_{\nu}^{\mathrm{L}}u_{\beta}(p_{\beta}),$ (3.4)
where $\gamma_{\mu}^{\mathrm{L}}=\gamma_{\mu}(1-\gamma^{5})/2$.
@@ 4.2 Dressed propagators of Dirac neutrinos in matter
$S_{1}^{-1}-GS_{2}G$ $=\Sigma_{11}^{-1},\quad S_{2}^{-1}-GS_{1}G=\Sigma_{22}^{-1}.$ (4.7)
@@
$S_{a}^{-1}$ $=-\frac{4(p_{0}-E_{a-}-g_{a}/2+\mathrm{i}0)}{(1-\alpha_{a}^{2})(1-\xi^{2})}\left(\begin{array}[]{cc}0&1+\xi(\bm{\sigma}\hat{\mathbf{p}})\\ 1+\xi(\bm{\sigma}\hat{\mathbf{p}})&0\end{array}\right)(1+\alpha_{a}\gamma^{5}),$ (4.11)
where we use the notation $\hat{\mathbf{p}}=\mathbf{p}/p$ for a unit vector along $\mathbf{p}$.
@@
$D_{1}^{-1}$ $=\frac{g(1+\alpha)/2-A(a-b)(1+\xi^{2})+2A\xi(a-b)(\bm{\sigma}\hat{\mathbf{p}})}{\left[g(1+\alpha)/2-A(a-b)(1+\xi^{2})\right]^{2}-4A^{2}\xi^{2}(a-b)^{2}},$
$D_{2}^{-1}$ $=\frac{g(1-\alpha)/2-A(a+b)(1+\xi^{2})+2A\xi(a+b)(\bm{\sigma}\hat{\mathbf{p}})}{\left[g(1-\alpha)/2-A(a+b)(1+\xi^{2})\right]^{2}-4A^{2}\xi^{2}(a+b)^{2}}.$ (4.15)
@@
$D_{1}^{-1}\to-\frac{g[1-(\bm{\sigma}\hat{\mathbf{p}})]}{2[(p_{0}-E_{1-}-g_{1}/2+\mathrm{i}0)(p_{0}-E_{2-}-g_{2}/2+\mathrm{i}0)-g^{2}]},$ (4.16)
@@
$\Sigma_{12}=-\frac{g}{2[(p_{0}-E_{1-}-g_{1}/2+\mathrm{i}0)(p_{0}-E_{2-}-g_{2}/2+\mathrm{i}0)-g^{2}]}\left(\begin{array}[]{cc}0&0\\ 1-(\bm{\sigma}\hat{\mathbf{p}})&0\end{array}\right),$ (4.17)
@@
$\Sigma_{11}$ $=-\frac{p_{0}-E_{2-}-g_{2}/2}{2[(p_{0}-E_{1-}-g_{1}/2+\mathrm{i}0)(p_{0}-E_{2-}-g_{2}/2+\mathrm{i}0)-g^{2}]}\left(\begin{array}[]{cc}0&0\\ 1-(\bm{\sigma}\hat{\mathbf{p}})&0\end{array}\right),$ (4.20)
$\Sigma_{22}$ $=-\frac{p_{0}-E_{1-}-g_{1}/2}{2[(p_{0}-E_{1-}-g_{1}/2+\mathrm{i}0)(p_{0}-E_{2-}-g_{2}/2+\mathrm{i}0)-g^{2}]}\left(\begin{array}[]{cc}0&0\\ 1-(\bm{\sigma}\hat{\mathbf{p}})&0\end{array}\right).$ (4.23)
@@
$S_{a}\to-\frac{1}{2(p_{0}-E_{a-}-g_{a}/2+\mathrm{i}0)}\left(\begin{array}[]{cc}0&0\\ 1-(\bm{\sigma}\hat{\mathbf{p}})&0\end{array}\right).$ (4.24)
@@ 4.3 Matrix element and transition probability in matter
To proceed with the calculation of the matrix element in Eq. (3) we choose the coordinate system so that $\mathbf{L}=L\mathbf{e}_{z}$, with $L>0$. Then, for simplicity, we assume the forward scattering approximation for charged leptons. In this case, $u_{e,\mu}^{\mathrm{T}}=(0,0,0,1)$. Thus, both the incoming electron and the outgoing muon are left particles propagating along the $z$-axis. Moreover, we rewrite the energies of left neutrinos as $E_{a-}\approx E_{a}+g_{a}/2$, where $E_{a}=\sqrt{q^{2}+m_{a}^{2}}$ is the energy of a free particle. We also use the cylindrical coordinates $(\rho,z,\phi)$ for the momentum of a virtual neutrino, $\mathbf{q}=\rho\mathbf{e}_{\rho}+z\mathbf{e}_{z}$.
@@
$\mathcal{M}_{e\to\mu}\approx-\frac{\mathrm{i}\bar{E}_{m}e^{\mathrm{i}\bar{E}_{m}L}}{2\pi L\mathfrak{E}_{m}}\sin(\mathfrak{E}_{m}L)\left[\left(\frac{\Delta m^{2}}{4E}+\frac{g_{2}-g_{1}}{2}\right)\sin 2\theta+g\cos 2\theta\right]\left(1+\frac{\mathrm{i}}{2L\bar{E}_{m}}\right),$ (4.27)
@@
and $\Delta m^{2}=m_{2}^{2}-m_{1}^{2}>0$. We neglect terms $\sim(\mathfrak{E}_{m}/\bar{E}_{m})^{2}\ll 1$ in deriving Eq. (4.27).
@@
with help of Eqs. (2.2) and (4.4). Taking into account Eqs. (4.2) and (4.27)-(4.29), one gets the transition probability for flavor oscillations of Dirac neutrinos in matter, $P_{\nu_{e}\to\nu_{\mu}}\propto|\mathcal{M}_{e\to\mu}|^{2}$.
In squaring of the matrix element, we omit the term $E^{2}_{m}/4\pi^{2}\approx E^{2}/4\pi^{2}$ since it is the common multiplier in the transition probability. In practice, this factor can be eliminated by the proper normalization of the interaction rate of incoming charged leptons with a nucleus in a neutrino source. Additionally, one has the factor $1/L^{2}$ which the transition probability is proportional to. It stays to provide the inverse square law which the flux of neutrinos, emitted in a source, obeys. Since this factor is the same for all neutrino types, i.e. it does not influence the dynamics of neutrino oscillations, we also omit it.
Eventually, we get that the total transition probability can be represented in the following form: $P_{\nu_{e}\to\nu_{\mu}}=P_{\nu_{e}\to\nu_{\mu}}^{(\text{qm})}+P_{\nu_{e}\to\nu_{\mu}}^{(\text{qft})}$ where
$P_{\nu_{e}\to\nu_{\mu}}^{(\text{qm})}=$ $\frac{\left(\frac{\Delta m^{2}}{4E}\sin 2\theta\right)^{2}}{\left(\frac{\Delta m^{2}}{4E}\sin 2\theta\right)^{2}+\left(\frac{\Delta m^{2}}{4E}\cos 2\theta-\frac{G_{\mathrm{F}}n_{e}}{\sqrt{2}}\right)^{2}}$
$\times\sin^{2}\left(\sqrt{\left(\frac{\Delta m^{2}}{4E}\sin 2\theta\right)^{2}+\left(\frac{\Delta m^{2}}{4E}\cos 2\theta-\frac{G_{\mathrm{F}}n_{e}}{\sqrt{2}}\right)^{2}}L\right),$
$P_{\nu_{e}\to\nu_{\mu}}^{(\text{qft})}=$ $P_{\nu_{e}\to\nu_{\mu}}^{(\text{qm})}/4L^{2}E^{2}.$ (4.30)
In Eq. (4.3), we keep the terms up to $L^{-2}$.
Formally, $P_{\nu_{e}\to\nu_{\mu}}^{(\text{qm})}$ in Eq. (4.3) coincides with the prediction of the QM approach for neutrino oscillations in matter [2, 3]. However, in QM, the quantity $E$ is referred to a mean energy of a neutrino beam, $E_{\nu}$, which is a not well defined parameter if one deals with a system of flavor neutrinos. We recall that, in the QFT treatment, $E=(E_{\alpha}+E_{\beta})/2$. Thus, this quantity is related to the energies of charged leptons, which are the observables since these particles correspond to in and out states.
The correction to the transition probability, given by $P_{\nu_{e}\to\nu_{\mu}}^{(\text{qft})}$, results from the QFT treatment of neutrino oscillations. This quantity is suppressed when a sufficiently long propagation distance is considered. However, if $L$ is macroscopic but relatively short, then $P_{\nu_{e}\to\nu_{\mu}}^{(\text{qft})}$ can contribute to the total probability.
Analogous corrections to the probability were studied in Ref. [22], where it was suggested that this kind of contributions can resolve the reactor antineutrino anomaly (see, e.g., Ref. [23]). Note that the computed correction to the transition probability turns out to be positive, $P_{\nu_{e}\to\nu_{\mu}}^{(\text{qft})}=\delta>0$. Hence, the total survival probability is suppressed by the same amount, $P_{\nu_{e}\to\nu_{e}}=P_{\nu_{e}\to\nu_{e}}^{(\text{qm})}-\delta$. Therefore, our result is consistent with Ref. [22].
@@ 5.1 Interaction of Dirac neutrinos with an electromagnetic field
where we use the fact that the mixing matrix in Eq. (2.2) is unitary.
If only the magnetic field is present, $\mathbf{E}=0$ and $\mathbf{B}\neq 0$, we shall see shortly in Sec. 5.2 that the transitions between neutrinos belonging to different flavors and opposite helicity states are possible provided that these neutrinos have a transition magnetic moment. This process is called the spin-flavor precession.
@@ 5.2 QM description of the neutrino spin-flavor precession
As in Sec. 4, we study two flavor neutrinos, $\nu_{e}$ and $\nu_{\mu}$. As a rule, neutrino oscillations are described in the flavor basis in frames of the QM approach. However, owing to the fact that both the nature of neutrinos and the neutrino magnetic moments are defined for neutrino mass eigenstates, we adopt the mass basis here. The corresponding mass eigenstates are supposed to be Dirac particles. We take that these massive neutrinos have only the transition magnetic moment $\mu$ and interact with a magnetic field $\mathbf{B}=(B,0,0)$ transverse with respect to the neutrino momentum which is along the $z$-axis.
@@ 5.3 Dressed propagators of Dirac neutrinos in magnetic field
$\mathbf{l}$ $=\hat{\mathbf{p}}\left(\frac{2\varsigma_{1}(p_{0}-E_{1}+\mathrm{i}0)}{1-\varsigma_{1}^{2}}+\frac{\mu^{2}B^{2}\varsigma_{2}}{2(p_{0}-E_{2}+\mathrm{i}0)}\right)-\frac{\mu^{2}\varsigma_{2}(\mathbf{B}\hat{\mathbf{p}})\mathbf{B}}{(p_{0}-E_{2}+\mathrm{i}0)}.$ (5.31)
@@
$\Sigma_{11}=\frac{(p_{0}-E_{2})(\gamma^{0}-\bm{\gamma}\hat{\mathbf{p}})}{2[(p_{0}-E_{1}+\mathrm{i}0)(p_{0}-E_{2}+\mathrm{i}0)-\mu^{2}B^{2}+\mu^{2}(\mathbf{B}\hat{\mathbf{p}})^{2}]}.$ (5.32)
@@
$\Sigma_{22}=\frac{(p_{0}-E_{1})(\gamma^{0}-\bm{\gamma}\hat{\mathbf{p}})}{2[(p_{0}-E_{1}+\mathrm{i}0)(p_{0}-E_{2}+\mathrm{i}0)-\mu^{2}B^{2}+\mu^{2}(\mathbf{B}\hat{\mathbf{p}})^{2}]}.$ (5.33)
@@
The denominators of $\Sigma_{aa}$ in Eqs. (5.32) and (5.33) contain the term $\propto\mu^{2}(\mathbf{B}\cdot\hat{\mathbf{p}})^{2}$. Analogous terms are present in the dressed propagators of Majorana neutrinos in a magnetic field, which were obtained in Ref. [18]. In the QM description of the spin-flavor precession in Sec. 5.2, the magnetic field is taken to be transverse to the neutrino momentum. Thus, in the QM approach, one is allowed to set $(\mathbf{B}\cdot\hat{\mathbf{p}})=0$ since neutrinos are on the mass shell. In QFT treatment, we cannot drop the term $\propto\mu^{2}(\mathbf{B}\cdot\hat{\mathbf{p}})^{2}\neq 0$ in Eqs. (5.32) and (5.33) since $\mathbf{p}$ is the momentum of a virtual neutrino. In fact, $\mathbf{p}$ can be arbitrary since we integrate over it in Eq. (3). Note that the quantity in question, $\propto\mu^{2}(\mathbf{B}\cdot\hat{\mathbf{p}})^{2}$, is an essentially quantum term in the propagator which arises from the summation of the infinite number of the Feynman diagrams in Fig. 2. Shortly in Sec. 5.4, we shall compute the contributions to the matrix element and the transition probability originating from this term.